# Perseus: Removing Energy Bloat from Large Model Training

### v1

@@ 3.1 Optimization Approach
By controlling the execution speed of individual computations in the pipeline, our goal is to minimize the energy consumption of a pipeline, and we can safely slow down a pipeline’s iteration time until the straggler’s iteration time $T^{\prime}$:
$\min_{F}$ $\quad\textrm{Energy}(F)$ (1)
@@
where $F$ is the set of GPU frequencies to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively. Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ on the “iteration time–energy” 2D plane. However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier, a monotonically decreasing tradeoff curve.
Figure 3 depicts the three cases where the straggler’s iteration time $T^{\prime}$ can possibly be w.r.t. the “iteration time–energy” Pareto frontier. $T_{min}$ is the shortest iteration time on the frontier, which coincides with the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption. Together, $T_{\min}$ and $T^{*}$ bookend the Pareto frontier.
  1. 1.
    When there are no stragglers (Figure 2(a)), we simply operate on the point on the Pareto frontier with the same iteration time, which reduces intrinsic energy bloat.
  2. 2.
    When a moderately slow straggler is detected (Figure 2(b)), we additionally reduce extrinsic energy bloat by slowing down non-straggler pipelines until $T^{\prime}$, using up all the slack time created by the straggler.
  3. 3.
    Finally, when the straggler’s iteration time goes beyond the minimum-energy point $T^{*}$ on the frontier (Figure 2(c)), we stop at $T^{*}$ instead of $T^{\prime}$. This is because slowing down beyond $T^{*}$ will instead increase energy.
@@
$T_{\textrm{opt}}=\min(T^{*},T^{\prime}).$ (2)
Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier, we know the solution of Equation 1 for any feasible value of $T^{\prime}$. Then our system will be able to react quickly to emerging stragglers by simply looking up the set of frequencies $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$. We present our algorithm to efficiently characterize the entire “iteration time–energy” Pareto frontier in Section 4.
3.2 Perseus Architecture
@@
Perseus represents each iteration of the training pipeline as a static directed acyclic graph (DAG), where nodes are forward and backward computations and edges are dependencies in between computations. Each node on the computation DAG is annotated with its planned time and energy consumption, which we call the energy schedule. Perseus realizes an energy schedule by executing each computation with a specific GPU frequency.
@@
Figure 4 shows the high-level architecture of Perseus. Perseus is split into a framework-independent server and a framework-integrated client. The server runs our time–energy optimization algorithm, produces all energy schedules on the “iteration time–energy” Pareto frontier, and caches them in a table for fast lookup. The client profiles computations of the pipelines online during training, and realizes the energy schedule by changing the GPU’s frequency during runtime.
Figure 4: Perseus architecture and workflow.
@@
In Perseus, a training job is specified by its computation DAG. When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy consumption of each forward and backward computation in each stage while training is running, and sends the profile to the server.
As training proceeds, 2 the server asynchronously characterizes the “iteration time–energy” Pareto frontier of the training job (§4), after which 3 the Pareto-optimal energy schedule corresponding to the shortest possible iteration time $T_{\min}$ is deployed to the client. Energy schedules are realized by the client’s Asynchronous Frequency Controller, integrated into the training framework’s pipeline execution engine (§5). Pareto-optimal energy schedules are saved in a lookup table format index by $T^{\prime}$.
During training, 4 the hardware infrastructure (e.g., data center rack power manager) notifies the Perseus server of a straggler pipeline and how much slowdown it will experience. The Perseus server can 5 quickly react to this by looking up the Pareto-optimal energy schedule corresponding to that straggler iteration time, and deploying it to the clients.
4 Algorithm Design
In this section, we describe our algorithm to obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail. We start by formulating the problem (§4.1). Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3). Finally, we extend our algorithm to support tensor parallelism, single-choice operations, and diverse pipeline schedules (§4.4).
4.1 Problem Formulation
@@
The energy consumption of a pipeline is not only from computation; it is the sum of three parts: (1) Computation; (2) Blocking on communication between computations; and (3) Blocking on communication until the straggler pipeline finishes. Formally,
$\sum_{i}{e_{i}(f_{i})}+P_{\textrm{blocking}}(N\cdot T-\sum_{i}{t_{i}(f_{i})})+P_{\textrm{blocking}}\cdot N\cdot(T^{\prime}-T)$ (3)
@@
Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline depends on the straggler’s iteration time $T^{\prime}$ in 2.222This makes sense because if the straggler slows down more and more, the non-straggler will consume more energy waiting for the straggler to finish. Specifically, the Pareto frontier for any $T^{\prime}$ is the Pareto frontier of $\textrm{Time}(F)$ vs. 1 shifted upwards by 2. Here, 1 does not depend on $T^{\prime}$. Therefore, if we characterize the Pareto frontier of $\textrm{Time}(F)$ vs. 1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$ as shifting the frontier upwards does not affect iteration time nor the frequency assignments that lead to each iteration time. Thus, we define
$\textrm{Energy}(F)=\sum_{i}{\left(e_{i}(f_{i})-P_{\textrm{blocking}}\cdot t_{i}(f_{i})\right)}$ (4)
@@
Finding one point on the “iteration time–energy” Pareto frontier with iteration time $T^{\prime}$ is equivalent to solving the following optimization problem:
$\min_{F}$ $\quad\textrm{Energy}(F)$ (5)
s.t. $\quad\textrm{Time}(F)\leq T^{\prime}$
@@
Reduction from 0/1 Knapsack. Details in Appendix B. ∎
The complete Pareto frontier can be obtained by solving PEM for all $T^{\prime}\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is in turn because the possible choices of GPU frequencies are discrete. However, if choices were continuous, the problem is exactly and efficiently solvable [57]. This is akin to integer linear programs becoming tractable when relaxed to linear programs.
We fit the exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements to make the choices continuous. We choose the exponential function due to its inherent flexibility and natural fit to data. We show in Section 6 that this relaxation produces high-quality approximate solutions that lead to actual savings. After relaxation, the problem’s optimization variables are the time and energy consumption planned for each computation in the pipeline, or the energy schedule.
4.2 Algorithm Overview
Iteratively Discovering the Pareto Frontier.
Figure 5: Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
While the relaxed Pipeline Energy Minimization problem is no longer NP-Hard, solving the problem for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient. Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate the next energy schedule adjacent to it on the Pareto frontier? Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, and every energy schedule we encounter will be Pareto-optimal.
@@ 4.2 Algorithm Overview
We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) is incorrect. That energy schedule is not Pareto-optimal because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
TL;DR.
@@
5 $p\leftarrow$ Minimum energy for all computations $\mathcal{P}\leftarrow\{p\}$ while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
@@
6       $p\leftarrow$ ReduceTime($\mathcal{G}$, $p$, $\tau$) $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
@@
4.3 Reducing Time with Minimum Energy Increase
Figure 6: A simplified example of how to reduce iteration time by unit time $\tau$. Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2). Then the Critical DAG ( 3) is obtained by considering every and only the computations on the critical path. Our key observation is that any valid s-t cut on the Critical DAG will reduce the iteration time by unit time $\tau$. Cut A and Cut B are two examples of valid s-t cut ( 4). Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
In this section, we describe our core subroutine ReduceTime (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
@@ 4.3 Reducing Time with Minimum Energy Increase
Originally, Perseus’s representation of the computation DAG is node-centric, which has forward and backward computations as nodes and their dependencies as edges. As a setup for subsequent steps, we convert this into an edge-centric computation DAG where computations are edges and dependencies are nodes (i.e., all incoming edges must complete before any outgoing edge can begin). This conversion can be done by splitting each node on the original computation DAG into two nodes and connecting the two with an edge annotated with the computation on the original node.
@@
Our goal is to reduce the execution time of the computation DAG by $\tau$, which is equivalent to reducing the length of all critical paths by $\tau$. Since computations that are not on any critical path (i.e., non-critical computations) do not affect the length of the critical path, we simply remove them from the computation DAG.
@@
Figure 6 shows two examples of this. We have already 1 transformed an energy schedule into 2 a computation DAG where computations are represented by edges, and 3 removed all non-critical computations. Here, 4 shows two valid s-t cuts: Cut A and Cut B. 5A speeds up the Backward computation cut by Cut A from $3\tau$ to $2\tau$, and the iteration time of the energy schedule was reduced by $\tau$. Similarly, 5B speeds up the Forward computation and Backward computation defined by Cut B from $5\tau$ to $4\tau$ and from $7\tau$ to $6\tau$ respectively, and the iteration time of the energy schedule was also reduced by $\tau$. Especially, in the second case, iteration time was only reduced because computations on two parallel critical paths were sped up together.
@@
We have shown with examples that any s-t cut represents a way to reduce the duration of the computation DAG by $\tau$. Then, a natural question is, which cut brings the smallest possible energy increase?
We can precisely map the capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of the exponential function fit for that computation and defining it to be the edge’s flow capacity. Then, our problem reduces to minimum cut, which we can solve with maximum flow. Appendix D has further details.
Converting to GPU Frequencies.
After finding the minimum capacity (minimum energy increase) cut, we will modify the durations of the computations involved in the cut, which results in a new energy schedule. Finally, we will convert the energy schedule into GPU frequencies that can be realized by the Perseus client. For each computation $i$, we convert its execution time $t$ to the slowest GPU frequency that will execute faster than $t$. This is because when computations are tightly packed by our algorithm, while slightly speeding up a computation is acceptable, slowing down any computation on the critical path will directly slow down the entire DAG, increasing intrinsic energy bloat.
@@
Our optimization algorithm has polynomial runtime. Let $N$ and $M$ respectively denote the number of stages and microbatches. Then, the computation DAG will have $O(NM)$ number of nodes and edges, and maximum flow with Edmonds-Karp runs in $O(N^{3}M^{3})$. While for general DAGs the total number of steps is known to be exponential to the size of the DAG [58], we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$. See Appendix E for proof.
In reality, commonly used number of stages ($N$) are 4 to 8 as too many increase pipeline bubble ratio [43, 14]. The number of microbatches ($M$) is typically around $4N$ [26, 59], but recently with high data parallel degree, far less have been reported even for high-performance settings [14]. As such, the algorithm runtime is negligible in practical scenarios (§6.5), given that large model training time is typically several weeks [48].
@@
Input: DAG $\mathcal{G}$ with computations $i$
Current energy schedule $p$
@@
Output: Pareto-optimal schedule with reduced time $p^{\prime}$
3
@@
       $\mathcal{G}\leftarrow$ Split nodes into two and connect with edge $\mathcal{G}\leftarrow$ Crash bipartite node set to one node $\triangleright$ Remove non-critical computations ( 3)
5       Annotate earliest & latest start times for $\forall i\in\mathcal{G}$ for $i\in\mathcal{G}$ do
             $\triangleright$ Critical computations should have zero slack

### v2

@@ 3.1. Unified Optimization Framework
Our goal is to minimize a pipeline’s energy consumption by controlling the execution speed of each computation in the pipeline. In doing so, we can slow down a pipeline’s iteration time at most until the straggler’s iteration time $T^{\prime}$:
(1) $\min_{F}$ $\quad\textrm{Energy}(F)$
@@
where $F$ is the set of GPU frequencies [^2] The SM frequency of NVIDIA GPUs can be set via NVML (nvml, ) in around 10 ms (AMD GPUs via AMDSMI (amdsmi, )), which is much shorter than typical large model computation latencies. Locking the GPU’s frequency provides deterministic computation latency (nvml, ; clockwork-osdi20, ), making it suitable for tightly planning and packing execution over time. to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively. Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ on the “iteration time–energy” 2D plane. However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier.
Now, let us assume we have a fully characterized “iteration time–energy” Pareto frontier with $T_{\min}$ and $T^{*}$ bookending the frontier (described in Section 4). $T_{\min}$ is the shortest iteration time on the frontier, which is the same as the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption. Figure 3 depicts the three cases where the straggler’s iteration time $T^{\prime}$ can possibly be w.r.t. the Pareto frontier:
  1. (1)
    When there are no stragglers (Figure 3(a)), we simply operate on the point on the Pareto frontier with iteration time $T_{\min}$, which reduces intrinsic energy bloat.
  2. (2)
    When a moderately slow straggler is detected (Figure 3(b)), we additionally reduce extrinsic energy bloat while keeping intrinsic bloat low by slowing down all non-straggler pipelines until $T^{\prime}$, using up all the slack time.
  3. (3)
    Finally, the straggler’s iteration time may go beyond the minimum-energy point $T^{*}$ on the frontier (Figure 3(c)). In this case, we only slow down non-stragglers until $T^{*}$, because going past $T^{*}$ will instead increase energy.
@@
(2) $T_{\textrm{opt}}=\min(T^{*},T^{\prime}).$
Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier by finding all Pareto-optimal $F$s, Perseus will be able to react instantly to emerging stragglers by looking up $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$.
Finally, we note that unlike other problem settings that do not consider energy consumption, fully utilizing all the slack time created by the straggler is not always energy-optimal. Being too fast or too slow can both waste energy.
3.2. Perseus Architecture
@@
Perseus represents each iteration of the training pipeline as a static directed acyclic graph (DAG), where nodes are forward and backward computations in each stage and edges are dependencies between computations. Each node on the computation DAG is annotated with its planned time and energy consumption, which we call the energy schedule. Perseus realizes an energy schedule by executing each computation with a specific GPU frequency.
@@
Figure 4 shows the high-level architecture of Perseus. Perseus is split into a framework- and hardware-agnostic server and a framework-integrated and device-specific client. The server pre-characterizes the “iteration time–energy” Pareto frontier (§4) and caches them for fast lookup. The client profiles pipeline computations online during training and realizes the energy schedule by changing the GPU’s frequency during runtime (§5).
Figure 4. Perseus architecture and workflow.
@@
In Perseus, a training job is primarily specified by its computation DAG for one training iteration. When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy of each forward and backward computation on each frequency while training is running.
Upon receiving profile results, 2 the server characterizes the “iteration time–energy” Pareto frontier (§4), and Pareto-optimal energy schedules are saved in a lookup table indexed by $T^{\prime}$. Then, 3 the energy schedule corresponding to the shortest iteration time is deployed to the client. Energy schedules are realized by the client’s Asynchronous Frequency Controller, integrated into the training framework (§5).
During training, 4 the training infrastructure (e.g., datacenter rack power/temperature manager) notifies the Perseus server of a straggler and its degree of slowdown. The server then 5 quickly reacts to this by looking up the Pareto-optimal energy schedule corresponding to the anticipated straggler iteration time, and deploys it to the client.
4. Characterizing the Pareto Frontier
In this section, we describe our algorithm to efficiently obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail. We first formulate the problem, show that it is NP-hard, and provide a relaxation (§4.1). Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3). Finally, we extend our algorithm to support 3D/hybrid parallelism, constant-time operations, and diverse pipeline schedules (§4.4).
4.1. Problem Formulation
@@
The energy consumption of a pipeline is not only from computation; it is the sum of three parts: (1) Computation; (2) Blocking on communication between computations; and (3) Blocking on communication until the straggler pipeline finishes:
(3) $\sum_{i}{e_{i}(f_{i})}+P_{\textrm{blocking}}(N\cdot T-\sum_{i}{t_{i}(f_{i})})+P_{\textrm{blocking}}\cdot N\cdot(T^{\prime}-T)$
@@
As derived in Section 3.1, given straggler iteration time $T^{\prime}$, we draw a vertical line on the “iteration time–energy” Pareto frontier ($\textrm{Time}(F)$ vs. 1+ 2) at $T_{\textrm{opt}}$ and find $F_{\textrm{opt}}$ where the two lines intersect. Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline only depends on the straggler’s iteration time $T^{\prime}$ in 2, which is only a constant upwards shift of the frontier. Therefore, if we characterize the Pareto frontier of $\textrm{Time}(F)$ vs. 1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$. Thus, we define
(4) $\textrm{Energy}(F)=\sum_{i}{\left(e_{i}(f_{i})-P_{\textrm{blocking}}\cdot t_{i}(f_{i})\right)}$
@@
Finding one point on the Pareto frontier with iteration time $T$ is equivalent to solving the following optimization problem:
(5) $\min_{F}$ $\quad\textrm{Energy}(F)$
s.t. $\quad\textrm{Time}(F)\leq T$
@@
Reduction from Knapsack. Details in Appendix C. ∎
The complete Pareto frontier can be obtained by solving PEM for all $T\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is because the possible choices of GPU frequencies are discrete. However, if choices were continuous, the problem is exactly and efficiently solvable (skutella1998, ). This is akin to integer linear programs becoming tractable when relaxed to linear programs.
The transform from the original problem to the relaxed version is done by fitting a continuous exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements for each forward and backward computation. We show in Section 6.2 that this relaxation produces high-quality approximate solutions that realizes most of the opportunity for savings. Solving the relaxed problem returns the time and energy consumption planned for each computation in the pipeline, or the energy schedule. Then, this is transformed back to a feasible solution of the original problem, which is the set of GPU frequencies $F$ (Section 4.3).
4.2. Iteratively Discovering the Pareto Frontier
Figure 5. Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
Although our relaxed problem is no longer NP-hard, solving it for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient. Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate its neighbor energy schedule on the Pareto frontier? Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, discovering every Pareto-optimal energy schedule.
@@ 4.2. Iteratively Discovering the Pareto Frontier
We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) would be incorrect. That energy schedule is not Pareto-optimal because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
@@
7 $p\leftarrow$ Minimum energy for all computations
8 $\mathcal{P}\leftarrow\{p\}$
9
10while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
@@ 8 $\mathcal{P}\leftarrow\{p\}$
11       $p\leftarrow$ GetNextPareto($\mathcal{G}$, $p$, $\tau$)
12       $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
13
@@ 12       $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
4.3. Finding the Neighbor Pareto-Optimal Schedule
Figure 6. A simplified example of how to reduce iteration time by unit time $\tau$. Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2). Then the Critical DAG ( 3) is obtained by considering every and only the computations on the critical path. Our key observation is that any valid s-t cut on the Critical DAG will reduce the iteration time by unit time $\tau$. Cut A and Cut B are two examples of valid s-t cut ( 4). Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
In this section, we describe our core subroutine GetNextPareto (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
@@ 4.3. Finding the Neighbor Pareto-Optimal Schedule
Originally, Perseus’s representation of the computation DAG is node-centric, which has forward and backward computations as nodes and their dependencies as edges. As a setup for subsequent steps, we convert this into an edge-centric computation DAG where computations are edges and dependencies are nodes (i.e., all incoming edges must complete before any outgoing edge can begin). This conversion can be done by splitting each node into two and connecting the two with an edge annotated with the computation on the original node.
@@
Our goal is to reduce the execution time of the computation DAG by $\tau$, which is equivalent to reducing the length of all critical paths by $\tau$ [^5] Let’s say there are two critical paths that run in parallel. They must be of equal length to both be critical paths. Here, if only one were shortened, the other will remain the sole critical path and the DAG will not execute faster. Since computations that are not on any critical path (i.e., non-critical computations) do not affect the length of the critical path, we remove them from the computation DAG.
@@
Figure 6 shows two examples of this. 4 shows two valid s-t cuts: Cut A and Cut B. 5A speeds up the computation edge cut by Cut A from $3\tau$ to $2\tau$, and the iteration time of the energy schedule was reduced by $\tau$. Similarly, 5B speeds up the computation edges cut by Cut B from $5\tau$ to $4\tau$ and from $7\tau$ to $6\tau$, and the iteration time of the energy schedule was also reduced by $\tau$. Especially, in the second case, iteration time was only reduced because computations on two parallel critical paths were sped up together.
@@
We have shown that an s-t cut represents a way to speed up the entire DAG by $\tau$. But speeding up computations increases energy. Then, a natural question is, which cut brings the smallest energy increase?
We can precisely map the flow capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of its exponential function (§4.1) and defining it to be the edge’s flow capacity. Then, our problem reduces to minimum cut, which we can solve with maximum flow. After finding the minimum cut, we modify the durations of the computations involved in the cut, obtaining a new Pareto-optimal energy schedule. Appendix D provides details with mathematical expressions for flow capacities.
Converting Back to GPU Frequencies.
Finally, we convert the energy schedule into GPU frequencies that can be realized by the Perseus client. For each computation, we convert its planned execution time $t$ to the slowest GPU frequency that will execute faster than $t$. This is because when computations are tightly packed by our algorithm, while slightly speeding up a computation is acceptable, slowing down any computation on the critical path will directly slow down the entire DAG, increasing intrinsic energy bloat.
API Description
profiler.begin(type) Begin time and energy profiling for computation type.
profiler.end(type) Record time and energy profiling results for computation type.
controller.set_speed(type) Set the hardware’s execution speed as planned for computation type.
server.set_straggler(id, delay, degree) Notify that a straggler is anticipated after delay seconds. A straggler returning to normal can be communicated by setting degree to 1.
Table 2. The minimal set of Perseus client and server APIs that require implementation. One client process manages each accelerator. The type parameter should be either "forward" or "backward". On GPUs, the “speed” control knob is the SM frequency. The server API will issue an HTTP/RPC request to the server. The id parameter is the ID of an accelerator.
@@
Our optimization algorithm has polynomial runtime. Let $N$ and $M$ denote the number of stages and microbatches, respectively. Then, the computation DAG will have $O(NM)$ number of nodes and edges, and maximum flow with Edmonds-Karp runs in $O(N^{3}M^{3})$. While for general DAGs the total number of steps is known to be exponential to the size of the DAG (skutella-thesis, ), we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$. See Appendix E for proof.
In reality, commonly used number of stages ($N$) is 4 to 8 (at most tens) to reduce pipeline bubble ratio (megatronlm-sc21, ; mlperf_training_31, ). Number of microbatches ($M$) is typically around $4N$ (gpipe-neurips19, ; megatron-turing-arxiv22, ), but recently with high data parallel degree, far less have been reported even for high-performance settings (mlperf_training_31, ). As such, algorithm runtime is practically negligible (§6.5), especially given that large model training easily takes weeks or months (patterson2021carbon, ).
@@
3
Input: DAG $\mathcal{G}$ of computations $i\in\mathcal{G}$
Current Pareto-optimal energy schedule $p$
@@
Output: Pareto-optimal schedule w/ reduced time $p^{\prime}$
@@
7 $\mathcal{G}\leftarrow$ Split nodes into two and connect with edge
8
$\triangleright$ Find and remove non-critical computations ( 3)
9 Annotate earliest & latest start times for $\forall i\in\mathcal{G}$
10 for $i\in\mathcal{G}$ do